OpenStudy (anonymous):

solve the following system of equations. {10x+4y=24 -6-4y=8 Part 1: combine the two equations using the elimination method. what is then value of x?

OpenStudy (anonymous):

Sorry I have no Idea. I wish I could help.

OpenStudy (anonymous):

thanks anyways !

OpenStudy (anonymous):

24 - 6 -4y = 8 Can you solve this?

OpenStudy (anonymous):

@m_galvan

OpenStudy (anonymous):

Can you solve for y in the equation 26 -6 -4y = 8?

OpenStudy (anonymous):

not sure how u did that

OpenStudy (anonymous):

You are supposed to solve one equation at a time

OpenStudy (anonymous):

would you start with 10x+4y=24?

OpenStudy (anonymous):

NO because it would be easier to solve for ONE variable FIRST. Do you know what variables are?

OpenStudy (anonymous):

x and y?

OpenStudy (anonymous):

Yes. X and Y are variables. To solve for a variable, you have to isolate it by moving all the other numbers to the other side (of the equal sign). To move the numbers to the other side (you need to move each number one by one), you have to change its sign to the opposite sign - for example x + 8 = 10 becomes x + 8 -8 = 10 -8. You have to this to both sides to isolate x. Do you get it?

OpenStudy (anonymous):

yes I get that

OpenStudy (anonymous):

So you know how to solve for x?

OpenStudy (anonymous):

yes I do I think I got it from here now thanks so much!